OpenStudy (anonymous):

y=x^3+2x^2-4x-8 list the intercepts

OpenStudy (anonymous):

you know the y-intercept right?

OpenStudy (anonymous):

ummm nope.. im guesin -8??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yeahhhhhhhh but the for the make y=o

OpenStudy (anonymous):

(-2, 0)(2, 0)

OpenStudy (anonymous):

okay so i got my y intercept. how do i find my x intercept?

OpenStudy (anonymous):

Solve for x, like I did above.

OpenStudy (anonymous):

one factor is x+2

OpenStudy (anonymous):

oh my g this confusing!!!!

OpenStudy (dumbcow):

use factor by grouping, it works out nicely (x^3 +2x^2) + (-4x-8) =0 x^2(x+2) -4(x+2) =0 (x+2)(x^2-4) = 0 x+2 = 0 x = -2 x^2 -4 = 0 x^2 = 4 x = +-2

OpenStudy (anonymous):

my points are (2,0), (-2,0) and (0,-8)?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

thank you